OpenStudy (anonymous):

I need help/answer! Medal will be given. Write the equation in slope-intercept form of the line that has a slope of 5 and y-intercept of -3

OpenStudy (amistre64):

how do you define the slope intercept form?

OpenStudy (anonymous):

That I have no idea. I think its something like y=mx+b

OpenStudy (amistre64):

thats actually correct

OpenStudy (amistre64):

now, m is a placeholder for the slope; and b is a placeholder for the value of the y intercept just plug and play

OpenStudy (anonymous):

y=5x-3?

OpenStudy (amistre64):

excellent

OpenStudy (anonymous):

Now I have another problem I would like help on.

OpenStudy (anonymous):

*Write the equation in slope-intercept form of the line that has a slope of -3 and contains the point (4,-5)

OpenStudy (anonymous):

use y-y1=m(x-x1) form then subtract the value of y1 from both sides

OpenStudy (anonymous):

I can not get the answer for that.

OpenStudy (anonymous):

All I have so far is -3x+12

OpenStudy (anonymous):

subtract 5 from the 12 so you would get -3x+7

OpenStudy (anonymous):

But my y1 is a -5

OpenStudy (anonymous):

yeah the form is y-(-5) which equals to y+5, so u subtract five from both side to leave the y by itself

OpenStudy (anonymous):

So the answer is y=-3x+7?

OpenStudy (anonymous):

Yes